OpenStudy (ttp):

What is the approximate distance from (2, 5) to (-2, 1)? 6 units 5 units 4 units 3 units

OpenStudy (anonymous):

use the distance formula

OpenStudy (anonymous):

http://0.tqn.com/d/math/1/0/A/1/distance.gif

OpenStudy (gorv):

root(2-(-2))^2+(5-1)^2)=root(4^2+4^2)=4*root(2)

OpenStudy (anonymous):

so, \[\sqrt{(2-(-2))^2+(5-1)^2}\]\[\sqrt{32}\]\[4\sqrt{2}\]

OpenStudy (anonymous):

The answer is 4.